OpenStudy (cookiemonster18):

what is the slope of the line that passes through the points (-2,5) and (1,4) A.-3 B.-2 C.-1/3 D.1/3

OpenStudy (alexandervonhumboldt2):

use slope formula: (y1-y2)/(x1-x2)=substitute coordinates: (5-4)/(-2-1)=1/(-3)=-1/3